# Tensor Product Attention Is All You Need

### v4

@@ 5.1 Language Modeling Tasks
All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy2022, ), using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (touvron2023llama, ) with SwiGLU activation (shazeer2020glu, ) and RoPE embeddings (su2024roformer, ), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention ) with Multi-Query Attention (MQA; shazeer2019fast ), Grouped Query Attention (GQA; ainslie2023gqa ), or Multi-head Latent Attention (MLA; liu2024deepseek ). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). Details on architecture hyperparameters and training hardware are shown in Appendix J.1.
@@
Training & Validation Curves. Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Training loss curves are provided in Appendix Figure 3. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix I) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses. For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek, ) (blue curves) generally trains more slowly and yields higher validation losses.
@@
Validation Perplexity. Figure 9 (in the Appendix) shows the validation perplexities of the medium- and large-scale models. Mirroring the loss curves, TPA and TPA-KVonly steadily outperform MHA, MQA, GQA, and MLA over the course of training. By the end of pretraining (around $49$B tokens), TPA-based approaches achieve the lowest perplexities in most configurations.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC, ), BoolQ (BoolQ, ), HellaSwag (hellaswag, ), OBQA (OpenBookQA, ), PIQA (PIQA, ), WinoGrande (WinoGrande, ), and MMLU (MMLU, ), using the lm-evaluation-harness codebase (eval-harness, ). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 12 and 13 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 141516 and 17 in the appendix.
@@
For the medium-size (353M) models (Table 2 for 0-shot and Table 15 in appendix for 2-shot), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%. When given two-shot prompts, TPA again leads with 53.12% average accuracy. A similar trend appears for the large-size (773M) models (Table 3), where TPA-KVonly attains the highest average (53.52% zero-shot). For the XL size models (1.5B) (Table 13 in the appendix), TPA-KV only achieves the highest average (55.03% zero-shot).
Our experiments confirm that TPA consistently matches or exceeds the performance of established attention mechanisms (MHA, MQA, GQA, MLA) across medium and large model scales.
@@ 5.2 Experimental Results on FlashTPA Decoding
The evaluations were performed with batch sizes selected from $\{1,2,4,8,16\}$, model embedding dimensions ($d_{\text{model}}$) chosen from $\{1024,2048,3072\}$, and sequence lengths ranging from $2^{12}$ (4,096) to $2^{19}$ (524,288). For all experiments, the dimension per head ($d_{h}$) was fixed at 64. The ranks for TPA’s factorized components ($R_{Q},R_{K},R_{V}$) were set to $(16,1,1)$, and for GQA configurations, the number of key-value head groups was 4.
The decoding time per token, measured as $\log_{2}(\text{time})$ in seconds, is plotted against
$\log_{2}(\text{sequence length})$. Lower values on the y-axis indicate faster decoding times. Results are presented in Figure 5 for an embedding dimension of 2048 (corresponding to 32 attention heads). Additional results for embedding dimensions of 1024 (16 heads, Figure 8) and 3072 (48 heads, Figure 7) are provided in Appendix D. Figure 5 depicts these speed comparisons for an embedding dimension of 2048. The results indicate that FlashTPA (blue line) is highly competitive and often outperforms other attention mechanisms, especially as the sequence length increases. A detailed breakdown of these comparisons across different batch sizes for the $d_{\text{model}}=2048$ case (Figure 5) is provided in Appendix D.
These findings underscore the computational efficiency of the FlashTPA decoding algorithm, particularly its favorable scaling with increasing sequence lengths.
@@ 6 Conclusion
Appendix A Expressing MHA, MQA, and GQA as Non-contextual TPA
This appendix section details how standard Multi-Head Attention (MHA), Multi-Query Attention (MQA), and Grouped-Query Attention (GQA) can be expressed as special, non-contextual variants of Tensor Product Attention (TPA).
A.1 MHA as Non-contextual TPA
Standard multi-head attention (MHA) can be viewed as a specific instance of TPA in which: 1) the rank is set equal to the number of heads; 2) the head dimension factor is non-contextual (i.e., independent of the $t$-th token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$); 3) the token dimension factor is a linear function of $\mathbf{x}_{t}$. More precisely, if we set $R_{Q}=R_{K}=R_{V}=h$ (the number of heads) and define the factors appropriately, MHA emerges. To match MHA with TPA, let $R_{Q}=R_{K}=R_{V}=h$. Focusing on $\mathbf{Q}_{t}$:
  1. (a)
    Non-contextual head factors. Define
    $\mathbf{a}^{Q}_{i}=h\mathbf{e}_{i}\in\mathbb{R}^{h},$ (A.1)
    where $\mathbf{e}_{i}\in\mathbb{R}^{h}$ is the $i$-th standard basis vector, so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
  2. (b)
    Contextual token factors. Define
    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$ (A.2)
    where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the per-head query projection matrix, hence $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$ is dependent on $\mathbf{x}_{t}$.
@@ A.1 MHA as Non-contextual TPA
$\mathbf{Q}_{t}=\sum_{i=1}^{h}\Big{[}\mathbf{e}_{i}\otimes\bigl{(}(\bm{W}^{Q}_{i})^{\top}\,\mathbf{x}_{t}\bigr{)}\Big{]}\in\mathbb{R}^{h\times d_{h}}.$ (A.3)

### v5

@@ 6.1 Language Modeling Tasks
All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022, ), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (touvron2023llama, ) with SwiGLU activation (shazeer2020glu, ) and RoPE embeddings (su2024roformer, ), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention ) with Multi-Query Attention (MQA; shazeer2019fast ), Grouped Query Attention (GQA; ainslie2023gqa ), or Multi-head Latent Attention (MLA; liu2024deepseek ). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). We pretrain all models with 50B tokens using 1 epoch. Details on architecture hyperparameters and training hardware are shown in Appendix H.1.
@@
Training & Validation Curves. Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Training loss curves are provided in Appendix Figure 3. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix G) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses. For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek, ) (blue curves) generally trains more slowly and yields higher validation losses.
@@
Validation Perplexity. Figure 9 (in the Appendix) shows the validation perplexities of the medium- and large-scale models. Mirroring the loss curves, TPA and TPA-KVonly steadily outperform MHA, MQA, GQA, and MLA over the course of training. By the end of pretraining (around 4949B tokens), TPA-based approaches achieve the lowest perplexities in most configurations.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC, ), BoolQ (BoolQ, ), HellaSwag (hellaswag, ), OBQA (OpenBookQA, ), PIQA (PIQA, ), WinoGrande (WinoGrande, ), and MMLU (MMLU, ), using the lm-evaluation-harness codebase (eval-harness, ). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 11 and 12 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 131415 and 16 in the appendix.
@@
For the medium-size (353M) models (Table 2 for 0-shot and Table 14 in appendix for 2-shot), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%. When given two-shot prompts, TPA again leads with 53.12% average accuracy. A similar trend appears for the large-size (773M) models (Table 3), where TPA-KVonly attains the highest average (53.52% zero-shot). For the XL size models (1.5B) (Table 12 in the appendix), TPA-KV only achieves the highest average (55.03% zero-shot). Our experiments confirm that TPA consistently matches or exceeds the performance of established attention mechanisms (MHA, MQA, GQA, MLA) across medium and large model scales.
@@ 6.2 Experimental Results on FlashTPA Decoding
The evaluations were performed with batch sizes selected from $\{1,2,4,8,16\}$, model embedding dimensions ($d_{\text{model}}$) chosen from $\{1024,2048,3072\}$, and sequence lengths ranging from $2^{12}$ (4,096) to $2^{19}$ (524,288). For all experiments, the dimension per head ($d_{h}$) was fixed at 64. The ranks for TPA’s factorized components ($R_{Q},R_{K},R_{V}$) were set to $(16,1,1)$, and for GQA configurations, the number of key-value head groups was 4. The decoding time per token, measured as $\log_{2}(\text{time})$ in seconds, is plotted against $\log_{2}(\text{sequence length})$. Lower values on the y-axis indicate faster decoding times. Results are presented in Figure 5 for an embedding dimension of 2048 (corresponding to 32 attention heads). Additional results for embedding dimensions of 1024 (16 heads, Figure 8) and 3072 (48 heads, Figure 7) are provided in Appendix B. Figure 5 depicts these speed comparisons for an embedding dimension of 2048. The results indicate that FlashTPA (blue line) is highly competitive and often outperforms other attention mechanisms, especially as the sequence length increases.
@@ 7 Conclusion
Acknowledgements
We thank the anonymous reviewers and area chair for their helpful comments. We acknowledge the compute credits provided by Fetch.ai.
Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
Our objective in this section is to compute attention without explicitly forming $\mathbf{Q},\mathbf{K},\mathbf{V}$, by contracting their factorized representations in a cache- and throughput-friendly order. Recall from Section˜3.1 that each per-token slice $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ is a sum of rank-$1$ outer products. Unless otherwise stated we use the per-factor normalizations $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$.
We make the batch/time/head/rank/value dimensions explicit and introduce the shorthands $D:=d_{h}$ and $E:=d_{v}$ (typically E=DE{=}D):
$\mathbf{A}_{Q}\in\mathbb{R}^{B\times T_{q}\times h\times R_{Q}},\quad\mathbf{B}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times D},\qquad\mathbf{A}_{K}\in\mathbb{R}^{B\times T_{k}\times h\times R_{K}},\quad\mathbf{B}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times D},$
$\mathbf{A}_{V}\in\mathbb{R}^{B\times T_{k}\times h\times R_{V}},\qquad\mathbf{B}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times E}.$
Indices $b,q,k,h,r,s,u,d,e$ denote batch, query position, key position, head, query-rank, key-rank, value-rank, feature ($D$), and value feature ($E$). We write $T:=T_{q}\!=\!T_{k}$ for full-sequence attention; in decoding, $T_{q}\!=\!1$ and we denote the cache length by $M\!=\!T_{k}$.
High-level idea. We first compute head-shared feature-space dot products between $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$, then mix them with head-specific $\mathbf{A}_{Q},\mathbf{A}_{K}$ to obtain logits, apply the masked softmax, and finally aggregate values via $\mathbf{A}_{V},\mathbf{B}_{V}$. This ordering avoids materializing any $T_{q}\times h\times D$ queries/keys/values.
@@ Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
Figure 6: Specialized TPA computation without materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$. Phase 1 (top): compute head-shared feature-space dot products $P[b,q,k,r,s]{=}\langle\mathbf{B}_{Q}[b,q,r,:],\,\mathbf{B}_{K}[b,k,s,:]\rangle$ and mix them with head-specific factors $\mathbf{A}_{Q},\mathbf{A}_{K}$ to obtain logits $\mathcal{L}[b,h,q,k]$. Phase 2 (bottom): apply the causal/padding mask and softmax to get $\alpha[b,h,q,k]$, then aggregate values via $\mathbf{A}_{V},\mathbf{B}_{V}$. Scalings $s_{Q},s_{K},s_{V}$ and $1/\sqrt{D}$ are folded into the corresponding phases. Dropout is omitted for clarity. Batch $B$, heads $H$, ranks $R_{Q},R_{K},R_{V}$, and feature dims $D,E$ are indicated in the nodes.
A.1 Direct computation in factor space
Single head. For a fixed head $h\!\in\![H]$ and token indices $(q,k)$, using $s_{Q}{=}1/R_{Q}$ and $s_{K}{=}1/R_{K}$ we have
$\bigl[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}\bigr]_{q,k}\;=\;\frac{1}{R_{Q}R_{K}}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\,a^{(r)}_{q,h}(\mathbf{x}_{q})\,a^{(s)}_{k,h}(\mathbf{x}_{k})\;\big\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{q}),\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{k})\big\rangle,$ (A.1)
and for values (with $s_{V}{=}1/R_{V}$), $\mathbf{V}^{(h)}_{k}=\frac{1}{R_{V}}\sum_{u=1}^{R_{V}}a^{(u)}_{k,h}(\mathbf{x}_{k})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{k})$. The per-head attention output at query position $q$ is then $\sum\nolimits_{k}\operatorname{softmax}\!\big(\frac{1}{\sqrt{D}}[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}]_{q,:}\big)_{k}\,\mathbf{V}^{(h)}_{k}$.
Multi-head with head-shared feature dot-products. Define head-shared feature-space dot products $P[b,q,k,r,s]\!=\!\langle\mathbf{B}_{Q}[b,q,r,:],\mathbf{B}_{K}[b,k,s,:]\rangle$. With $\mathcal{S}\!=\!\frac{s_{Q}s_{K}}{\sqrt{D}}$, we compute
$\mathcal{L}[b,h,q,k]$ $=\mathcal{S}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\mathbf{A}_{Q}[b,q,h,r]\;\mathbf{A}_{K}[b,k,h,s]\;P[b,q,k,r,s],$ (A.2)
$\alpha[b,h,q,k]$ $=\operatorname{Softmax}_{k}\!\left(\mathcal{L}[b,h,q,k]+\log\mathsf{mask}[q,k]\right),$
$\mathbf{O}[b,h,q,e]$ $=s_{V}\sum_{k=1}^{T_{k}}\sum_{u=1}^{R_{V}}\alpha[b,h,q,k]\;\mathbf{A}_{V}[b,k,h,u]\;\mathbf{B}_{V}[b,k,u,e].$ (A.3)
Here $\mathsf{mask}[q,k]\!\in\!\{0,-\infty\}$ enforces causality and padding. Eqs. (A.2)–(A.3) make explicit that (i) feature-space dot products $P$ are head-shared, and (ii) all $1/R_{(\cdot)}$ scalings can be folded into a single multiplicative factor per phase without affecting numerical stability (softmax shift invariance).
A.2 Complexity: materialized vs. specialized computation
We compare two execution strategies. (i) Naïve/materialized: form $\mathbf{Q},\mathbf{K},\mathbf{V}$ explicitly and call standard kernels. (ii) Specialized: compute via Eq. (A.2)–(A.3) using head-shared feature-dot products and per-head rank contractions.
Standard MHA (baseline).
Ignoring projections, full-sequence attention uses $\Theta(BHT^{2}D)$ FLOPs for scores and $\Theta(BHT^{2}D)$ for value aggregation, i.e., $\mathcal{F}_{\text{MHA}}\;=\;2\,\Theta(BHT^{2}D)$.
TPA (materialized).
Forming $\mathbf{Q},\mathbf{K},\mathbf{V}$ from factors costs $\Theta(BTHD(R_{Q}{+}R_{K}{+}R_{V}))$ after the linear projections; subsequent attention uses the same $2\,\Theta(BHT^{2}D)$ as MHA.
TPA (specialized).
Using Eqs. (A.2)–(A.3) and writing $T_{q}{=}T_{k}{=}T$, the dominant FLOPs are
$\underbrace{\Theta(B\,T^{2}\,R_{Q}R_{K}\,D)}_{\text{feature dots }P}\;+\;\underbrace{\Theta(B\,H\,T^{2}\,R_{Q}R_{K})}_{\text{per-head rank combine}}\;+\;\underbrace{\Theta(B\,H\,T^{2}\,R_{V}\,E)}_{\text{value aggregation}}.$
Compared to $\mathcal{F}_{\text{MHA}}$, the specialized path reduces FLOPs whenever
$R_{Q}R_{K}\,D\;+\;H\,R_{Q}R_{K}\;+\;H\,R_{V}\,E\;<\;2\,H\,D.$ (A.4)
Dividing by HDHD yields $(R_{Q}R_{K}/H)+(R_{Q}R_{K}/D)+R_{V}(E/D)<2$. For E=DE{=}D and small ranks (e.g., $R_{Q}{=}R_{K}{=}R_{V}{=}1$), the inequality holds for typical $H,D\!\geq\!2$ and the benefit grows with larger $H$ or $D$.
Memory traffic and peak working set.
For full-sequence attention the naive path streams $\mathbf{Q},\mathbf{K},\mathbf{V}$ of size $\Theta(BTHD)$ each. The specialized path streams factors only and needs the head-shared $P$ tiles of size $\Theta(B\,T_{q}^{\text{tile}}\,T_{k}^{\text{tile}}\,R_{Q}R_{K})$ plus per-head tiles for the rank combine/value aggregation. In decoding with cache length $M$, the factorized KV cache uses $(R_{K}{+}R_{V})(h{+}D)$ numbers per token (cf. Section 3.3), vs. $2hD$ for MHA; this reduction directly lowers memory bandwidth pressure.
Algorithm 1 Specialized TPA (no explicit $\mathbf{Q},\mathbf{K},\mathbf{V}$; causal)